A 5% solution of cane sugar (molar mass 342) is isotonic with 1% of a solution of an unknown solute. The molar mass of unknown solute in g/mol is ?

Grade:11

1 Answers

Harsh kothari
6 years ago
π1=π2n1/v1=n2/v25/(342×0.1)=1/(M×0.1)M=68.4gπ is osmotic pressuren is no of moleV is volumeBy given %by vol 5g in100ml